# AMNS: Attention-Weighted Selective Mask and Noise Label Suppression for Text-to-Image Person Retrieval thanks: †: Corresponding author

### v1

@@ Abstract
Text-to-image person retrieval aims to retrieve images of person given textual descriptions, and most methods implicitly assume that the training image-text pairs are correctly aligned, but in practice, under-correlated and false-correlated problems arise for image-text pairs due to poor image quality and mislabeling. Meanwhile, the random masking augmentation strategy may incorrectly discard semantic content resulting in the problem of generating noisy pairings between image lexical elements and text descriptions. To solve these two problems, we propose a new noise label suppression method and alleviate the problem generated by random mask through an attention-weighted selective mask strategy. In the proposed noise label suppression method, the effect of noise labels is suppressed by preventing the model from being overconfident by considering the inverse KL scatter loss, which is combined with the weight adjustment focus loss to further improve the model’s recognition ability on difficult samples. On the other hand, Attention-Weighted Selective Mask processes the raw image through the EMA version of the image encoder, retaining some of the tokens with strong semantic associations with the corresponding text descriptions in order to extract better features. Numerous experiments validate the effectiveness of our approach in terms of dealing with noisy problems. The code will be available soon at https://github.com/RunQing715/AMNS.git.
@@ I Introduction
Text-to-image person retrieval utilizes textual descriptions to aid in the retrieval of person images [1, 2, 3]. Although many effective retrieval methods have been proposed in the field[4, 5, 6, 7, 8, 9, 10], almost all of them implicitly assume that all input training pairs are correctly aligned. However, in the real world, noise is prevalent due to factors such as environmental disturbances, imprecise linguistic descriptions, mislabeling, or missing information, making it difficult for the idealized assumptions to hold. In addition, the random masking strategy[11, 12, 13] in the process of image enhancement may incorrectly discard semantic content, resulting in noise matching between image patches and text description, and ultimately lead to the degradation of retrieval performance. Therefore, effective noise suppression methods and mask strategies for images are crucial for improving the performance of model retrieval.
@@
In particular, some images are mislabeled, or the image quality is too low, but are still treated as positive examples for cross-modal learning, resulting in the model learning the wrong correspondences, a phenomenon known as Noisy Correspondence (NC). To solve this problem, there are two main approaches: sample selection[14, 15, 16, 17] and robust loss functions[18, 19, 20]. The former usually utilizes the memory effect of the deep neural networks[21] to progressively differentiate and filter noisy data, but this may require additional computational resources and carries the risk of mistakenly deleting valid data. Comparatively, the latter approach aims to develop noise-tolerance loss functions to improve the robustness of model training against NC.
@@
Therefore, this paper proposes a new cross-modal matching loss named image-text Bidirectional Similarity Distribution Matching (BSDM) loss, which effectively suppresses the negative impact of noisy labels by taking into account the bidirectional relationship between the predictive distribution and the true distribution, and is complemented by the introduction of a Weight Adjustment Focus (WAF) loss to enhance the model’s ability to handle difficult samples. In addition, to address the problem that random mask may destroy the semantic association between graphs and texts, this paper introduces an Attention-Weighted Selective Mask (AWM) strategy, which aims to optimize the feature extraction process and improve the retrieval performance by accurately retaining the image regions that are strongly correlated with the textual descriptions and reducing the noisy pairings. The main contributions and innovations of this paper are as follows:
@@
  • We investigate two new and prevalent problems in text-to-image person retrieval tasks, namely, the Noisy Correspondence (NC) and the noisy pairing problem, which may mislead the model to learn incorrect semantic information. To the best of our knowledge, this paper may be the first work to explore the noise pairing problem caused by the random masking strategy in person retrieval tasks.
@@
  • We present AMNS: attention-weighted selective mask and noise label suppression framework, by introducing Attention-Weighted Selective Mask (AWM) strategy to alleviate the noise pairing problem between graphs and texts caused by random masks, and by proposing Bidirectional Similarity Distribution Matching (BSDM) loss is proposed to effectively suppress the negative impact caused by noisy labels, and Weight Adjustment Focal (WAF) loss is proposed to be combined with BSDM loss, which further improves the model’s recognition ability on difficult samples.
  • Experiments show that using AWM strategy outperforms random masking strategy, and that the integration of the proposed method for suppressing noisy labels into other researchers’ methods achieves an overall improvement in performance. Meanwhile, both the mAP index and mINP index of AMNS are almost the best among the comparison methods, which fully proves the superiority of our proposed method and can be better used in real-life retrieval systems.
@@ II METHOD
The overall diagram of our proposed framework is shown in Fig. 1, which consists of an image encoder ${\mathbf{E_{i}}}$, a text encoder ${\mathbf{E_{t}}}$ and an EMA-based image encoder $\overline{\mathbf{E_{i}}}$. Given an image, crop it and input it to $\overline{\mathbf{E_{i}}}$, a mask image is obtained using an AWM strategy to highlight image regions relevant to the textual description and fed into ${\mathbf{E_{i}}}$ together with the raw image, enabling ${\mathbf{E_{i}}}$ to extract a more accurate representation of the features. The model makes $\overline{\mathbf{E_{i}}}$ filter out features with lower weights by self-supervised learning, and combines with the proposed BSDM, which considers the inverse KL scatter, to inhibit the noisy labelling problem by inversely constraining the true label distribution through the predictive distribution. Meanwhile, combining with the proposed WAF further enhances the model’s attention and learning of difficult samples by adjusting the weight coefficients. In addition, we adopt ID loss to aggregate feature representations of the same identity, which further improves the model retrieval performance.
@@ II-B Noise Label Suppression Methods
BSDM mitigates the effect of NC by introducing inverse KL scatter, but may weaken the model’s learning for difficult samples (image-text similarity scores that are low, but are actually positive example samples, or similarity scores that are high, but are actually negative example samples). Therefore, we consider the introduction of focal loss[25] and introduce proportional weight coefficients $\alpha$,$\beta$ in front of it to improve it, and by adjusting the coefficient factor $\gamma$ we are able to adjust the model’s attention to the misclassified samples. We refer to this loss as Weight Adjustment Focal Loss (WAF). The image-to-text WAF loss is as follows:
@@
By combining the two approaches of BSDM and WAF, the model not only suppresses the effect of noisy labels, but also enhances the learning of difficult samples, thus achieving better performance in noisy data environments.
@@ III-C Ablation experiments
Complementarity between BSDM and WAF: As can be seen from No.5 in Table III, the addition of WAF improves the Rank1 metric by 0.26%, and the mAP and mINP metrics by 0.31% and 0.45%, respectively, which fully demonstrates that WAF further improves the model’s ability to recognize difficult samples when it is complemented with BSDM.
@@ IV Conclusion
In this paper, we explore Noise Correspondence (NC) and the noise pairing problem caused by random masking strategy. We introduce Attention-Weighted Selective Mask to alleviate the noise matching problem caused by random masking strategy, and propose a novel method, that is, Bidirectional Similarity Distribution Matching method, combined with the improved focus loss, effectively deal with the NC, and achieve excellent performance. We have carried out extensive experiments, which fully prove the superiority of the proposed method in solving the above problems, and prove that our method can find all the correct matches better.

### v2

@@ Abstract
Text-to-image person retrieval aims to retrieve images of person given textual descriptions, and most methods implicitly assume that the training image-text pairs are correctly aligned, but in practice, under-correlated and false-correlated problems arise for image-text pairs due to poor image quality and mislabeling. Meanwhile, the random masking augmentation strategy may incorrectly discard semantic content resulting in the problem of generating noisy matching between image lexical elements and text descriptions. To solve these two problems, we propose a new noise label suppression method and alleviate the problem generated by random mask through an attention-weighted selective mask strategy. In the proposed noise label suppression method, the effect of noise labels is suppressed by preventing the model from being overconfident by considering the inverse KL scatter loss, which is combined with the weight adjustment focal loss to further improve the model’s recognition ability on hard samples. On the other hand, Attention-Weighted Selective Mask processes the raw image through the EMA version of the image encoder, retaining some of the tokens with strong semantic associations with the corresponding text descriptions in order to extract better features. Numerous experiments validate the effectiveness of our approach in terms of dealing with noisy problems. The code will be available soon at https://github.com/RunQing715/AMNS.git.
@@ I Introduction
Text-to-image person retrieval utilizes textual descriptions to aid in the retrieval of person images [1, 2, 3]. Although many effective retrieval methods have been proposed in the field[4, 5, 6, 7, 8, 9, 10], almost all of them implicitly assume that all input training pairs are correctly aligned. However, in the real world, noise is prevalent due to factors such as environmental disturbances, imprecise linguistic descriptions, mislabeling, or missing information, etc., making it difficult for the idealized assumptions to hold. In addition, the random masking strategy[11, 12, 13] in the process of image enhancement may incorrectly discard semantic content, resulting in noisy matching between image patches and text description, and ultimately lead to the degradation of retrieval performance. Therefore, effective noise suppression methods and image masking strategy are crucial for improving the performance of model retrieval.
@@
In particular, some images are mislabeled, or the image quality is too low, but are still treated as positive examples for cross-modal learning, resulting in the model learning the wrong correspondences, a phenomenon known as Noisy Correspondence (NC). To solve this problem, there are two main approaches: sample selection[14, 15, 16, 17] and robust loss functions[18, 19, 20]. The former usually utilizes the memory effect of the deep neural networks[21] to progressively differentiate and filter noisy data, but this may require additional computational resources and carry the risk of mistakenly deleting valid data. Comparatively, the latter approach aims to develop noise-tolerance loss functions to improve the robustness of model training against NC.
@@
Therefore, this paper proposes a new cross-modal matching loss named image-text Bidirectional Similarity Distribution Matching (BSDM) loss, which introduces the inverse KL scattering while retaining the forward KL scattering, so that the model does not blindly pursue a perfect match with all the pairs of positive samples when fitting the data, but rather maintains a certain degree of flexibility, thus effectively suppressing the negative impact of noise labeling. And it is complemented by the introduction of a Weight Adjustment Focal Loss (WAF) based to enhance the model’s ability to handle hard samples. In addition, to address the problem that random mask may destroy the semantic association between image patches and text description, this paper introduces an Attention-Weighted Selective Mask (AWM) strategy, which aims to optimize the feature extraction process and improve the retrieval performance by accurately retaining the image regions that are strongly correlated with the textual descriptions and reducing the noisy matching. The main contributions and innovations of this paper are as follows:
@@
  • We investigate two problems in text-to-image person retrieval tasks, namely the Noisy Correspondence (NC) and the noisy matching problem caused by random mask strategy. To our knowledge, this paper may be the first to explore the noisy matching problem caused by random mask strategies in pedestrian retrieval tasks.
@@
  • We present AMNS: attention-weighted selective mask and noise label suppression framework and introduce an Attention-Weighted Selective Mask (AWM) strategy to alleviate the noisy matching problem caused by random masks. It is proposed that Bidirectional Similarity Distribution Matching (BSDM) can effectively suppress the influence of noise labels, and combined with Weight Adjustment Focal Loss (WAF), The recognition ability of the model on difficult samples is improved.
  • Experiments show that AWM strategy is superior to random masking strategy. In addition, the noise suppression method proposed in this paper is applied to other models, and the performance is significantly improved. Among the AMNS methods, mAP index and mINP index are almost the best.
@@ II METHOD
The overall diagram of our proposed framework is shown in Fig. 1, which consists of an image encoder ${\mathbf{E_{i}}}$, a text encoder ${\mathbf{E_{t}}}$ and an EMA-based image encoder $\overline{\mathbf{E_{i}}}$. Given an image, crop it and input it to $\overline{\mathbf{E_{i}}}$, a mask image is obtained using an AWM strategy to highlight image regions relevant to the textual description and fed into ${\mathbf{E_{i}}}$ together with the raw image, enabling ${\mathbf{E_{i}}}$ to extract a more accurate representation of the features. The model makes $\overline{\mathbf{E_{i}}}$ filter out features with lower weights by self-supervised learning, and combines with the proposed BSDM, which considers the inverse KL scatter, to inhibit the noisy labelling problem by inversely constraining the true label distribution through the predictive distribution. Meanwhile, combining with the proposed WAF further enhances the model’s attention and learning of hard samples by adjusting the weight coefficients. In addition, we adopt ID loss to aggregate feature representations of the same identity, which further improves the model retrieval performance.
@@ II-B Noise Label Suppression Methods
BSDM mitigates the effect of NC by introducing inverse KL scatter, but may weaken the model’s learning for hard samples (image-text similarity scores that are low, but are actually positive example samples, or similarity scores that are high, but are actually negative example samples). Therefore, we consider the introduction of focal loss[25] and introduce proportional weight coefficients $\alpha$,$\beta$ in front of it to improve it, and by adjusting the coefficient factor $\gamma$ we are able to adjust the model’s attention to the misclassified samples. We refer to this loss as Weight Adjustment Focal Loss (WAF). The image-to-text WAF loss is as follows:
@@
By combining the two approaches of BSDM and WAF, the model not only suppresses the effect of noisy labels, but also enhances the learning of hard samples, thus achieving better performance in noisy data environments.
@@ III-C Ablation experiments
Complementarity between BSDM and WAF: As can be seen from No.5 in Table III, the addition of WAF improves the Rank1 metric by 0.26%, and the mAP and mINP metrics by 0.31% and 0.45%, respectively, which fully demonstrates that WAF further improves the model’s ability to recognize hard samples when it is complemented with BSDM.
@@ IV Conclusion
In this paper, we explore Noise Correspondence (NC) and the noisy matching problem caused by random masking strategy. We introduce Attention-Weighted Selective Mask to alleviate the noisy matching problem caused by random masking strategy, and propose a novel method, that is, Bidirectional Similarity Distribution Matching method, combined with the improved focal loss, effectively deal with the NC, and achieve excellent performance. We have carried out extensive experiments, which fully prove the superiority of the proposed method in solving the above problems, and prove that our method can find all the correct matches better.